You lower the temperature of a sample of liquid carbon disulfide from 94.1°C until its volume contracts by 0.491% of its initial value. What is the final temperature of the substance? The coefficient of volume expansion for carbon disulfide is 1.15 × 10-3 (C°)–1.
Answer #1

We know that,

V=V_{0}(1+\alpha \Delta T)

Where V is the final volume, V_{0} is the original volume, \alpha is the coefficient of volume expansion and \Delta T = (Final Temperature - Initial Temperature)

The given data shows that,

\frac{(V-V_{0})}{V_{0}}=-0.491/100

The negative sign indicated contraction

\frac{V}{V_{0}}=1- \frac{0.491}{100}

\frac{V}{V_{0}}=0.99509

1+\alpha \Delta T=0.99509

\alpha \Delta T=0.99509-1=-0.00491

Final Temperature - Initial Temperature = -0.00491/\alpha

Initial Temperature = 94.1 °C

\alpha = 1.15*10^-3 (°C^-1)

Final Temperature = \frac{-0.00491}{0.00115}+94.1 = 89.83 °C